OpenStudy (anonymous):

A Bit string is a sequence of zero or more bits. The length of this string is the number of bits in the string. Give a recursive definition for the set of bit strings { 0r1s0r | r, s ∈ ℕ } . Note the number of 0’s must be equal, but the number of 1’s may be different from the number of 0’s.

OpenStudy (freckles):

r,s are elements of the naturals or {0,1} ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am sorry. one second.

OpenStudy (anonymous):

{ 0^r1^s0^r | r, s ∈ ℕ }

OpenStudy (freckles):

Ok that actually makes more sense

OpenStudy (anonymous):

{ 0^r1^s0^r | r, s E N }

OpenStudy (anonymous):

{ 0^r1^s0^r | r, s is element of Natural numbers } Sorry I couldnt figure out how to use the symbol.

OpenStudy (anonymous):

\[{ 0^r1^s0^r | r, s \epsilon \mathbb{N} }\]

OpenStudy (freckles):

Why close this? I think I prefer it open so more people can jump because I'm actually still thinking .

OpenStudy (freckles):

Well let's just think about the sequence \[f(r,s)=0^r 1 ^s 0^r ; r,s \in \mathbb{N} \\ f(1,1)=010 \\ f(1,2)=0110 \\ f(2,1)=00100 \\ f(2,2)=001100 \\ f(3,1)=0001000 \\ f(3,2)=00011000 \\ f(3,3)=000111000 \\ f(1,3)=01110 \\ f(2,3)=0011100 \\ \] ... so we are to think of a recursive form for this...

OpenStudy (freckles):

Well this doesn't answer the question...but we can find the lenth \[|f(1,1)|=2(1)+1 \\ |f(1,2)|=2(1)+2 \\ |f(2,1)|=2(2)+1 \\ |f(2,2)|=2(2)+2 \\ |f(3,1)|=2(3)+1\\|f(3,2)|=2(3)+2\] I hope you see a pattern and kind find the length of something more general like: \[|f(m,n)| \text{ where } m,n \in \mathbb{N} \]